OpenStudy (anonymous):

Choose the equation that could be used to find two consecutive integers whose sum is 69. (4 points) n + (n + 2) = 69 n + (n + 1) = 69 n - (n + 1) = 69 n + 2n = 69

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (anonymous):

B is what i would pick. n+(n+1)=69 n+n+1=69 combine like terms 2n+1=69 get n on one side 2n=68 Isolate n n=34 now for the second number add 1 to that since you subtracted it for the first number the two numbers would be 34 and 35